OpenStudy (anonymous):

f(x)=(x-7)/(x+2), g(x)= (-2x-7)/(x-1) Can anyone show me how to find the F(g)(x)) and g(f)(x))? and show how they are inverses?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

F(g(x)) is found by treating the value of g(x) as x and inserting it in f(x) f(x)=(x-7)/(x+2) F(g(x)) = (g(x) - 7)/ ( g(x) + 2) Can you put the calue of g(x) & evaluate further from here

OpenStudy (anonymous):

ohhh okay wait I think I figured it out.. nevermind! Thanks though!:)

OpenStudy (anonymous):

yw:))) do ask if you need any help